40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 11 09:18:44 2010, Alberto, Update, General, Measurement running 
    Reply  Mon Jan 11 11:06:53 2010, Alberto, Update, ABSL, Measurement running 
Message ID: 2502     Entry time: Mon Jan 11 11:06:53 2010     In reply to: 2500
Author: Alberto 
Type: Update 
Category: ABSL 
Subject: Measurement running 

Quote:

I'm working on the AbsL experiment. A measurement which involved the PRC locked is running at the moment.

Please make sure of not interfering with the interferometer until it is done. Thank you.

 I'm done for the moement.

I realized that I need to take into account somehow the DC power from the photodiode. By now the measurement of the transmitted beat's power is affected by the total power circulating inside of the PRC and thus it depends on the cavity alignment.

I closed the laser shutter and turned down the flipping mirrors.

I'm planning to restart measuring by 2.30pm today. Till then the interferometer is available.

ELOG V3.1.3-